OpenStudy (praxer):

evaluate int_{0}^{1}(x^3+1) as limit of a sum

OpenStudy (praxer):

Please help me with this sum.

OpenStudy (praxer):

\[ evaluate \int\limits_{0}^{1} ( X^3+1) as limit of a \sum\]

OpenStudy (abb0t):

What?

OpenStudy (praxer):

evaluate this as the limit of a sum

OpenStudy (abb0t):

Is this linear algebra or calculus II?

OpenStudy (praxer):

calculus II

OpenStudy (abb0t):

\[\int\limits_{0}^{1}(x^3+1)dx\]

OpenStudy (praxer):

yes,

OpenStudy (abb0t):

Only cuz i'm really tired right now am I going to give you the answer: \(\sf \color{red}{\frac{x^4}{4} +x}\) now evaluate it from 0 < x < 1

OpenStudy (praxer):

ohk,

OpenStudy (abb0t):

You should know from the Fundamental Theorem of Calculus Pt. II: \[\int\limits_{a}^{b} f(x)dx= F(b)-F(a)\]

OpenStudy (praxer):

I know the theorem,

OpenStudy (abb0t):

It's area under a curve @Mimi_x3

OpenStudy (abb0t):

By definition, if f(x) is continuous on [a,b], then u divide the inteval into n subintervals of equal width, \(\Delta\)x, and from each interval choise a point, x\(_i\), and is defined by as: \[\int\limits_{a}^{b}f(x)dx = \lim_{n \rightarrow \infty} \sum_{i=1}^{n}f(x_i)\Delta x\]

OpenStudy (abb0t):

its the net area between a function and the x-axis.